OpenStudy (anonymous):

Mikalee, a fighter pilot, ejects from her burning plane. Once she is safely on the ground, she launches a flare to attract the attention of a rescue plane. The initial velocity of the flare is 1600 ft/sec. What is the maximum height of the flare? 50 ft 400 ft 4,000 ft 40,000 ft

OpenStudy (anonymous):

We can use the following equation from kinematics: \[v_f ^2 = v_i ^2 + 2ad\]Where v_f and v_i are the final and initial velocities, a is acceleration due to gravity, and d is the distance. We're solving for the distance. When the flare is at its max height, its velocity is zero. So: \[0=(1600 ft/s)^2 + 2(-32 f/s^2)d\]Now, solve for d. Can you try solving this on your own?

OpenStudy (anonymous):

is there a more simple equation?

OpenStudy (anonymous):

I believe this is as simple as you're going to get.

OpenStudy (anonymous):

ok, can you take me step by step ?

OpenStudy (anonymous):

Sure. We'll square the 1600 ft/s and multiply the 2 with the -32 ft/s^2 to get: \[0=2560000 ft^2 / s^2 - (64 ft/s^2)d\]Now add 64d to both sides: \[d(64 ft/s^2)=2560000 ft^2 / s^2\]Can you do the last step?

OpenStudy (anonymous):

the answer is 40000?

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

so if you were trying to find the maximum height of the flare would you ........

OpenStudy (anonymous):

The most direct method is to use the kinematics formula I gave you above.